# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ 3.1 Markovian and Reciprocal Processes on the Tree
Recall that the IMF procedure is motivated by the characterisation of the SB solution as the unique Markov measure that is a mixture of its bridges. Our first contribution is to extend this characterisation to the tree-based setting. We begin with the following definitions, akin to those in the standard setting.
@@
[…]
@@ 3.2 Iterative Markovian Fitting on the Tree
Theorem 3.1 suggests that an analogous Iterative Markovian Fitting procedure could be used to solve the dynamic Tree-SB problem; in this section, we show that this is indeed the case. The constructions and proofs proceed in much the same way as in Shi et al. (2023), but crucially rely on the following simple decomposition of the KL divergence according to the tree structure, which applies for measures in $\mathcal{M}_{\mathcal{T}}$ and $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$.
@@
We now state the following result, which shows convergence of the IMF iterates to the TreeSB solution. This resembles Theorem 8 in Shi et al. (2023), and the proof proceeds in largely the same way using the appropriate modifications and in particular leveraging the KL decomposition in Lemma 3.2. We defer the details to the Appendix.
@@
Under mild conditions, the $\mathrm{TreeSB}_{\mathrm{dyn}}$ solution $\mathbb{P}^{*}$ is the unique fixed point of the IMF iterates $\mathbb{P}^{n}$, and we have $\lim_{n\rightarrow\infty}{\mathrm{KL}}(\mathbb{P}^{n}\|\mathbb{P})=0$.
@@ 3.3 Implementation
$L_{u,v}=\begin{cases}\sum_{e\in\mathcal{E}(u)}\tfrac{1}{\sigma^{2}T^{(u,v)}}&\textrm{if $u=v$, where $\mathcal{E}(u)$ is set of edges incidentto $u$},\\ -\tfrac{1}{\sigma^{2}T^{(u,v)}},&\textrm{if $(u,v)\in\mathcal{E}$}\\ 0&\textrm{otherwise}.\end{cases}$
Using the formula for conditional multivariate Gaussians in terms of the precision matrix, we have $Y_{\mathcal{S}^{c}}|Y_{\mathcal{S}}=y_{\mathcal{S}}\sim\mathcal{N}(\tilde{\mu},\tilde{\Sigma}),$ using the block matrix expressions $\tilde{\mu}=-(L_{\mathcal{S}^{c}\mathcal{S}^{c}})^{-1}L_{\mathcal{S}^{c}\mathcal{S}}y_{\mathcal{S}}$ and $\tilde{\Sigma}=(L_{\mathcal{S}^{c}\mathcal{S}^{c}})^{-1}$. Based on the structure of the tree, we can therefore obtain a sample of the bridging process by: (1) calculating this conditional joint distribution ${\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$ over the unseen vertices, (2) sampling $Y_{\mathcal{S}^{c}|\mathcal{S}}$ according this distribution, and then (3) drawing Brownian bridges between $Y_{u}$ and $Y_{v}$ along each edge in the tree.
@@
The above construction allows us to construct samples from a reciprocal process $\Pi$. Recall from Definition 3.4 that the Markovian projection then proceeds by performing individual Markovian projections along each edge. We therefore maintain a separate neural network drift function $v_{\theta_{e}}$ for each edge $e=(u,v)$, each of which is trained according to the bridge-matching loss
@@
After training the vector fields, let ${\mathbb{M}}$ denote the path measure of the resulting Markov process on $\mathcal{T}$. Constructing the next reciprocal process requires samples from ${\mathbb{M}}$, which can be obtained by running the learned diffusions from the root node $r\in\mathcal{S}$ along the tree. Note that while in theory we have ${\mathbb{M}}_{i}=\mu_{i}$ for each $i\in\mathcal{S}$, errors will accumulate in practice. We therefore follow Shi et al. (2023) and learn both forward and backward diffusion processes along each edge, which give equivalent representations of the path measures via the corresponding time-reversals (see Appendix for details). This enables simulation along the tree from any vertex in $\mathcal{S}$ to obtain samples from the coupling, helping to mitigate errors accumulating in the marginals. Note also that all edges train independently, so can be run in parallel for faster computation.
@@
In keeping with previous naming conventions, we call the proposed algorithm Tree Diffusion Schrödinger Bridge Matching (TreeDSBM). The method is summarised in Algorithm 1 with full implementation details in the Appendix, and we provide an illustration of the two-step procedure in Figure 1. The algorithm can be initialised at any coupling $\Pi_{\mathcal{S}}^{0}$ over $\mathcal{S}$ with correct marginals; a standard choice would be the independent coupling ${\Pi_{\mathcal{S}}^{0}=\bigotimes_{i\in\mathcal{S}}\mu_{i}}$.
@@ 3.4 Connection to Fixed-Point Wasserstein Barycentre Algorithms
The unregularised Wasserstein-2 barycentre $\nu^{*}=\operatorname*{arg\,min}_{\nu}\sum_{i}\lambda_{i}W_{2}^{2}(\mu_{i},\nu)$ can be characterised (under mild assumptions) by the property $\sum_{i}\lambda_{i}T_{i}(x)=x$ for $x\in\textrm{supp}(\nu)$, where $T_{i}$ are the corresponding OT maps from $\nu$ to $\mu_{i}$ (Agueh and Carlier, 2011; Álvarez Esteban et al., 2016). This property has motivated fixed-point-based methods for barycentre computation (Álvarez Esteban et al., 2016; Korotin et al., 2022). Such methods take an initial candidate barycentre $\nu_{0}$, and at each iteration compute the OT maps $T_{i}$ to construct the next iterate $\nu_{i+1}=\bar{T}\#\nu_{i}$ using the map $\bar{T}(x)=\sum_{i}\lambda_{i}T_{i}(x)$.
In our setting, recall that a barycentre corresponds to a star-shaped tree with fixed marginals on the leaves. In this case, the bridges of ${\mathbb{Q}}$ are conditioned on each of the leaf vertices and the only unknown point is $Y_{0}$ at the centre vertex. The conditional distribution ${\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$ simplifies to
@@
Our method can therefore be viewed as a natural counterpart of fixed-point methods for barycentre computation, adapted to the case of flow-based entropic OT solvers.
3.5 Extensions for Improving Convergence Speed
Several recent works have introduced strategies to accelerate convergence of the IMF scheme. One approach is to warmstart the IMF iterations with an initial coupling $\Pi_{0}$ consisting of samples from a static $\varepsilon$-OT solver applied to minibatches (Tong et al. (2024b), see also Pooladian et al. (2023), Tong et al. (2024a)). Similarly, one could initialise the TreeIMF procedure using samples obtained from static mmOT solvers applied to minibatches. Another approach is the $\alpha$-DSBM algorithm of De Bortoli et al. (2024), which mitigates the iterative nature of IMF by instead considering a continuous flow of path measures to the SB solution. Extending this idea to the tree-structured setting could also provide similar benefits. We provide further details and discussion in the Appendix.
@@ 4 Experiments
We now empirically validate the performance of the proposed TreeDSBM algorithm. We first focus on comparisons against the TreeDSB algorithm from Noble et al. (2023), as this considers exactly the same problem setting, and we follow their experimental setups. We then compare against other continuous Wasserstein-2 barycentre solvers, and show that TreeDSBM is competitive with state-of-the-art approaches. We include comparisons against the methods WIN (Korotin et al., 2022), W2CB (Korotin et al., 2021), and the recent method NOTWB (Kolesov et al., 2024a). These methods have demonstrated strong empirical performance in their respective works, and are chosen here to be representative of different approaches in the literature—WIN is an iterative method inspired by Álvarez Esteban et al. (2016), W2CB is an Input Convex Neural Network-based approach (Amos et al., 2017; Makkuva et al., 2020), and NOTWB is based on recent Neural OT methodology (Korotin et al., 2023). We use the implementations in their publicly available code, to which we provide the links in Appendix F. Finally, we provide an example to demonstrate a potential limitation of our approach under certain settings. Full experimental details are included in the Appendix.
@@
Figure 2: Comparison of the learned barycentre for TreeDSBM (6 IMF iterations) against TreeDSB (50 IPF iterations) and WIN. TreeDSB and TreeDSBM samples are generated from each leaf vertex $k$, and for WIN we plot samples using the weighted-pushforward expression for the barycentre. Also displayed is a close approximation to the ground-truth, using the in-sample method of Cuturi and Doucet (2014).
Table 2: Sinkhorn divergence to the ‘ground-truth’ barycentre, for barycentre samples generated from each leaf vertex $k$. The Sinkhorn divergence is computed with entropy regularisation 0.01, using 5000 generated samples and 1500 points approximating the ground-truth.
@@
TreeDSBM (6 IMF) 1.16 1.13 1.18
TreeDSB (50 IPF) 2.35 4.04 2.35
WIN $(\sum_{i}\lambda_{i}T_{i})\#\nu$ 1.17
We first examine the performance of TreeDSBM in a low-dimensional synthetic example, using the experimental setup of Noble et al. (2023). We compute the $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$-barycentre of a moon, spiral, and circle dataset with $\varepsilon=0.1$ (recall $\sigma=\sqrt{\tfrac{\varepsilon}{2}}$). We compare TreeDSBM ran for 6 IMF iterations against TreeDSB ran for 50 IPF iterations (using checkpoints provided by Noble et al. (2023)). Note that this is a challenging problem—the lower-dimensional and discontinuous structures in the marginals mean that the barycentre has a complex and fragmented support, and the transport maps to the barycentre are highly discontinuous.
In Figure 2, we plot the obtained samples from both methods, and for comparison also display the barycentre obtained using the in-sample method from Cuturi and Doucet (2014) to give a close approximation to the ground-truth. We remark that the aim of this experiment is not to outperform in-sample methods; it is known that such approaches perform well in low dimensions, but do not scale well as dimension increases. To quantitatively assess performance, we report the Sinkhorn divergence (Genevay et al., 2018) relative to this ‘ground-truth’ in Table 2. The results show that TreeDSBM significantly improves over TreeDSB in this setting, approximating the barycentre to a high degree of accuracy and at a much lower computational cost (with good convergence after only a few IMF iterations). It is able to successfully capture the complex nature of the barycentre in this challenging example (previously suggested as a potential limitation of dynamic solvers in Noble et al. (2023)), and the barycentres generated from different vertices $k$ exhibit improved consistency. For details of the experimental setup, see the Appendix.
We additionally report results for the iterative WIN method (Korotin et al., 2022). We see that TreeDSBM performs competitively with this strong baseline for continuous Wasserstein-2 solvers. The results reported for WIN are for the combined map $(\sum_{i}\lambda_{i}T_{i})\#\nu$, as in our experiments the barycentre generator $\nu=G\#\rho$ was unable to fit the true barycentre accurately, nor were the maps $T_{i}^{-1}\#\mu_{i}$. Additionally, we applied the W2CB (Korotin et al., 2021) and NOTWB (Kolesov et al., 2024a) algorithms to this example, but were unable find hyperparameters to make the algorithms to converge to the correct solution. We hypothesise that the neural maps may struggle to model the discontinuous transports well, and also that this challenging example may result in difficult loss landscapes.
We found TreeDSBM to exhibit stable training despite this challenging problem setting, and to also be the fastest of the algorithms to converge to the solution. We provide a runtime analysis, along with further results and discussion, in the Appendix.
@@
We also compare performance of TreeDSB and TreeDSBM on a higher dimensional image dataset, computing the $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$-barycentre between MNIST digits 2, 4, and 6 (LeCun et al., 2010). In this setting, TreeDSB is reported in Noble et al. (2023) to exhibit training instability for low entropy regularisation $\varepsilon$, causing it to struggle to match the marginal measures at the vertices. In contrast, the IMF approach of TreeDSBM ensures matching of these marginals, thus allowing the use of much smaller regularisation values. In fact, using a large regularisation for TreeDSBM would limit sample quality, because of the noise added when sampling from ${\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$ in the reciprocal process construction. We therefore use $\varepsilon=0.01$ for TreeDSBM with 4 IMF iterations, and display a visual comparison with results reported in Noble et al. (2023) in Algorithm 1 (with more samples plotted in the Appendix). While it is difficult to validate the accuracy of such solutions, the obtained samples from TreeDSBM appear to display a greater resemblance to state-of-the-art barycentre methods for similar problems (see e.g. Kolesov et al. (2024a)), while again requiring significantly less computational cost than TreeDSB.
@@
The previous experiments have shown TreeDSBM to improve over its IPF counterpart TreeDSB. We now provide a more detailed comparison with current strongly-performing methods for continuous Wasserstein-2 barycentre estimation, by reporting results for standard experiments in the literature.
As a real-world applications experiment, we compare the algorithms in the subset posterior aggregation setting, considered in previous continuous barycentre works such as Staib et al. (2017); Li et al. (2020); Fan et al. (2021); Korotin et al. (2021). The barycentre of subset posteriors is known to be close to the true posterior (Srivastava et al., 2018), and as such it can be efficient to compute posteriors on only subsets of datasets before aggregating them using a barycentre algorithm.
We work with the experimental setup and dataset used in Korotin et al. (2021) (the same dataset was also used previously in Li et al. (2020); Fan et al. (2021)), which uses Poisson and negative-binomial regressions on a bike-rental dataset (Fanaee-T, 2013). The aim is to predict the hourly number of bike rentals using features including day of the week, weather conditions, and more. The dataset is 8-dimensional and is split into 5 distinct subsets each of size 100,000. The ‘ground-truth’ barycentre consists of 100,000 samples from the full dataset posterior.
Following the literature, we report the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric between the ‘ground-truth’ and the obtained samples in Table 3. We ran TreeDSBM for 4 IMF iterations with $\varepsilon=0.001$; for full hyperparameters details for all algorithms, see the Appendix. For methods that generate from each marginal, we report the average over generations from each marginal, and for WIN we report results for the barycentre generator. The $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric is defined as
where the Bures-Wasserstein metric is defined as $\mathrm{BW}_{2}^{2}(\nu,\tilde{\nu})=W_{2}^{2}(\mathcal{N}(m_{\nu},\Sigma_{\nu}),\mathcal{N}(m_{\tilde{\nu}},\Sigma_{\tilde{\nu}}))$ for the respective means and covariances of the distributions.
Table 3: $\mathrm{BW}_{2}^{2}\mathrm{-UVP},\%$ for different algorithms on the subset posterior aggregation experiment in Korotin et al. (2021), evaluated using 100,000 samples.
@@
$\downarrow$ Negative Binomial 0.009 0.024 0.018 0.013
From Table 3, we see that all methods perform strongly; it is difficult to conclusively say which performs best, given that we do not have perfect access to the ground truth barycentre. We see that TreeDSBM certainly performs competitively with these state-of-the-art approaches. Moreover, we found TreeDSBM to display fast training—both TreeDSBM and NOTWB obtained good results after only 5 minutes of training, while W2CB took approximately 10 minutes and WIN took around 45 minutes. This runtime is for training the edges of TreeDSBM sequentially, so could be sped up further by a factor of close to 5 if trained simultaneously. For more details regarding runtimes, see the Appendix.
Higher-dim Gaussian Experiments
We also report results for computing the barycentre of Gaussian distributions, in increasingly high dimensions. This is a standard experiment in the literature, because in this setting the ground-truth barycentre is a Gaussian whose parameters can be calculated accurately using a fixed point method (Álvarez Esteban et al., 2016).
We follow the experimental setup previously used in (Korotin et al., 2021, 2022; Kolesov et al., 2024a), in which 3 Gaussian distributions and its ground-truth ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre are randomly generated, for each dimension in $\{64,96,128\}$. We report results for $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ (which measures the quality of the overall generated barycentre) as described above, and additionally report the $\mathrm{L}^{2}\mathrm{-UVP}$ metric, which measures the quality of the individual maps to the barycentre and is defined for each marginal as
@@
$d=64$ 0.20 0.04 0.08 0.18
$\downarrow\mathrm{BW}_{2}^{2}\mathrm{-UVP},\%$ $d=96$ 0.30 0.07 0.10 0.17
$d=128$ 0.38 0.12 0.14 0.29
$d=64$ 0.96 0.17 0.10 1.22
$\downarrow L^{2}\mathrm{-UVP},\%$ $d=96$ 1.20 0.20 0.10 1.14
@@
[…]
4.1 Ave, celeba! Benchmark
Finally, we provide an example that illustrates a potential limitation of our approach. As previously discussed, it is difficult to evaluate performance of barycentre algorithms in high dimensions due to the lack of a ground-truth. To combat this, Korotin et al. (2022) proposed the Ave, celeba! barycenter benchmark, which consists of 3 distributions of transformed CelebA faces (Liu et al., 2015), for which the $(\tfrac{1}{4},\tfrac{1}{2},\tfrac{1}{4})$-Wasserstein-2 barycentre recovers the true CelebA dataset. We are not aware of any previous attempts to scale flow-based generative modelling approaches to this setting. We apply TreeDSBM for this dataset for 2 IMF iterations, started from an initial coupling obtained by training bridge matching to a Gaussian along each edge; for full implementation details, see the Appendix.
Figure 4: Samples from leaf nodes of the Ave, celeba! dataset transported to the barycentre, using TreeDSBM for 2 IMF iterations and $\sigma=0.01$.
While our approach scales to this problem, our obtained samples do not match the visual quality of state-of-the-art results reported in Kolesov et al. (2024a). Many samples are transported well, but some pick up unwanted artifacts. This suggests a limitation of TreeDSBM for this particular benchmark—the true transport maps in this specific setting are in fact simple (primarily just colour changes), but it is difficult for TreeDSBM to learn this because communication between the edges is infrequent and only occurs after each IMF iteration. We thus hypothesise that the flow-based and iterative nature of our algorithm may not be a suitable inductive bias for this benchmark; we add a more detailed discussion in the Appendix. Nevertheless, TreeDSBM is able to scale flow-based approaches to this challenging high-dimensional setting, and we anticipate that our results could be improved by utilising more advanced implementation techniques from the flow-matching literature (such as those in Kim et al. (2025)), and by modifying the neural architecture to exhibit inductive biases more suitable for this problem setting. We leave such investigations for future work.

### v2

@@ 3.1 Markovian and reciprocal processes on the tree
Recall that the IMF procedure is motivated by the characterisation of the SB as the unique Markov measure that is a mixture of its bridges. Our first contribution is to extend this characterisation to the tree-based setting. We begin with the following definitions, akin to those in the standard case.
@@
Under mild assumptions (in the Brownian case, namely that $\int\lVert x\rVert^{2}{\mathrm{d}}\mu_{i}(x)<\infty$ and ${\mathrm{H}}(\mu_{i})<\infty$ for each $i\in\mathcal{S}$), there exists a unique solution to the dynamic TreeSB problem ($\mathrm{TreeSB}_{\mathrm{dyn}}$). The solution is the unique process $\mathbb{P}$ that is both Markov and in $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$ with correct marginals $\mathbb{P}_{i}=\mu_{i}$ for $i\in\mathcal{S}$.
@@ 3.2 Iterative Markovian Fitting on the tree
Theorem 3.1 suggests that an analogous Iterative Markovian Fitting procedure could be used to solve the dynamic TreeSB problem; in this section, we show that this is indeed the case. The constructions and proofs proceed in much the same way as in [77], but crucially rely on the following simple decomposition of the KL divergence according to the tree structure, which applies for measures in $\mathcal{M}_{\mathcal{T}}$ and $\mathcal{R}_{\mathcal{S}}({\mathbb{Q}})$.
@@
We now state the following result, which shows convergence of the IMF iterates to the TreeSB solution. This resembles Theorem 8 in [77], and the proof proceeds in largely the same way using the appropriate modifications and in particular leveraging the KL decomposition in Lemma 3.2. We defer the details to Appendix B.
@@
Under mild conditions, the $\mathrm{TreeSB}_{\mathrm{dyn}}$ solution $\mathbb{P}^{*}$ is the unique fixed point of the TreeIMF iterates $\mathbb{P}^{n}$, and we have $\lim_{n\rightarrow\infty}{\mathrm{KL}}(\mathbb{P}^{n}\|\mathbb{P})=0$.
@@ 3.3 Implementation
[…]
@@
The above construction allows us to construct samples from a reciprocal process $\Pi$. Recall from Definition 3.4 that the Markovian projection then proceeds by performing individual Markovian projections along each edge. We therefore maintain a neurally-parameterised drift function $v_{\theta_{e}}$ for each edge $e=(u,v)$, each of which is trained according to the bridge-matching loss
@@
After training the vector fields, let $\mathbb{M}$} denote the path measure of the resulting Markov process on $\mathcal{T}$. Constructing the next reciprocal process requires samples from $\mathbb{M}$}, which can be obtained by running the learned diffusions from the root node $r\in\mathcal{S}$ along the tree. Note that while in theory we have ${\mathbb{M}}_{i}=\mu_{i}$ for each $i\in\mathcal{S}$, errors will accumulate in practice. We therefore follow [77] and learn both forward and backward diffusion processes along each edge, which give equivalent representations of the path measures via the corresponding time-reversals (see Appendix C for details). This enables simulation along the tree from any vertex in $\mathcal{S}$ to obtain samples from the coupling, helping to mitigate errors accumulating in the marginals. Note also that all edges train independently, so can be learned in parallel for faster computation.
@@
In keeping with previous naming conventions, we call the proposed algorithm Tree Diffusion Schrödinger Bridge Matching (TreeDSBM). The method is summarised in Algorithm 1 with full implementation details in Appendix C, and we provide an illustration of the two-step procedure in Figure 1. The algorithm can be initialised at any coupling $\Pi_{\mathcal{S}}^{0}$ over $\mathcal{S}$ with correct marginals; a standard choice would be the independent coupling $\Pi_{\mathcal{S}}^{0}=\bigotimes_{i\in\mathcal{S}}\mu_{i}$}. We provide a simplified and more explicit version of the algorithm when used for barycentre computation in Appendix C.
@@ 3.4 Connection to fixed-point Wasserstein barycentre algorithms
Recall that the unregularised Wasserstein-2 barycentre can be computed using the iterative fixed-point method, where each iteration proceeds by pushing-forward the current iterate through the map $\sum_{i}\lambda_{i}T_{i}$. In our setting, the barycentre problem corresponds to a star-shaped tree with fixed marginals on the leaves. In this case, the bridges of $\mathbb{Q}$} are conditioned on each of the leaf vertices and the only unknown point is $Y_{0}$ at the centre vertex, so the conditional distribution ${\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$ simplifies to
@@
Our method can therefore be viewed as a natural counterpart of fixed-point methods for barycentre computation, adapted to the case of flow-based entropic OT solvers. While a naive approach might consider using IMF to solve each OT sub-problem in the fixed-point scheme, resulting in nested iterations, our approach shows that the IMF and fixed-point iterations can in fact be elegantly combined into a single iterative procedure, along with a well-understood theoretical grounding. In particular, the expensive OT map computations required for the fixed-point procedure can instead be switched out for inexpensive bridge-matching procedures. Each iteration is therefore cheap, and empirically we found the TreeDSBM algorithm to retain the fast convergence property of the fixed-point procedure in terms of the number of iterations required.
@@ 4 Experiments
Figure 2: Comparison of the learned barycentre for TreeDSBM (6 IMF iterations) against TreeDSB (50 IPF iterations) and WIN. TreeDSB and TreeDSBM samples are generated from each leaf vertex $k$, and for WIN we plot samples using the weighted-pushforward expression for the barycentre. Also displayed is a close approximation to the ground-truth, using the in-sample method of [18].
@@
TreeDSBM
(6 IMF)
1.14
$\pm$ 0.07
1.05
$\pm$ 0.07
1.08
$\pm$ 0.11
TreeDSB
(50 IPF)
2.35 4.04 2.35
WIN
$(\sum_{i}\lambda_{i}T_{i})\#\nu$
1.17
Synthetic 2dd barycentre
We first examine the performance of TreeDSBM in a low-dimensional synthetic example, using the experimental setup of [64]. We compute the $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$-barycentre of a moon, spiral, and circle dataset with $\varepsilon=0.1$ (recall $\sigma=\sqrt{\tfrac{\varepsilon}{2}}$). We compare TreeDSBM ran for 6 IMF iterations against TreeDSB ran for 50 IPF iterations (using checkpoints provided by [64]). In Figure 2, we plot the obtained samples from both methods, and for comparison also display the barycentre obtained using the in-sample method from [18] to give a close approximation to the ground-truth. To quantitatively assess performance, we report the Sinkhorn divergence [35] relative to this ‘ground-truth’ in Figure 2. The results show that TreeDSBM significantly improves over TreeDSB in this setting, approximating the barycentre to a high degree of accuracy and at a much lower computational cost (with good convergence after only a few IMF iterations). It is able to successfully capture the complex nature of the barycentre in this challenging example (previously suggested as a potential limitation of dynamic solvers in [64]), and the barycentres generated from different vertices $k$ exhibit improved consistency. For details of the experimental setup, see Appendix D.1.
We additionally report results for the iterative WIN method [47]. We see that TreeDSBM performs competitively with this strong baseline for continuous Wasserstein-2 solvers. The results reported for WIN are for the combined map $(\sum_{i}\lambda_{i}T_{i})\#\nu$, as in our experiments the barycentre generator $\nu=G\#\rho$ was unable to fit the true barycentre accurately, nor were the maps $T_{i}^{-1}\#\mu_{i}$. Additionally, we applied the W2CB [49] and NOTWB [45] algorithms to this example, but were unable find hyperparameters to make the algorithms to converge to the correct solution. We hypothesise that the neural maps may struggle to model the discontinuous transports well, and also that this challenging example may result in difficult loss landscapes. We found TreeDSBM to exhibit stable training despite this challenging problem setting, and to also be the fastest of the algorithms to converge to the solution. We provide a runtime analysis, along with further results and discussion, in Appendix D.1.
@@
We also compare performance of TreeDSBM with TreeDSB on a higher dimensional image dataset, computing the $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$-barycentre between MNIST digits 2, 4, and 6 [52]. In this setting, TreeDSB is reported in [64] to exhibit training instability for low entropy regularisation $\varepsilon$, causing it to struggle to match the marginal measures at the vertices. In contrast, the IMF approach of TreeDSBM ensures matching of these marginals, thus allowing the use of much smaller regularisation values. In fact, using a large regularisation for TreeDSBM would limit sample quality, because of the noise added when sampling from ${\mathbb{Q}}_{\mathcal{S}^{c}|\mathcal{S}}$ in the reciprocal process construction. We therefore use $\varepsilon=0.02$ for TreeDSBM with 4 IMF iterations, and display a visual comparison with results reported in [64] in Figure 3. While it is difficult to validate the accuracy of such solutions, the obtained samples from TreeDSBM appear to display a greater resemblance to state-of-the-art barycentre methods for similar problems (see e.g. [45]), while again requiring significantly less computational cost than TreeDSB. We remark that the TreeDSBM samples in Figure 3(b) contain a small amount of noise, which is to be expected as we are solving for an entropy-regularised problem. In Appendix D.2 we also display samples with the noise reduced or removed, along with full experimental details and additional results.
@@
The previous experiments have shown TreeDSBM to improve over its IPF counterpart TreeDSB. We now provide a more detailed comparison with existing strongly-performing methods for continuous Wasserstein-2 barycentre estimation, namely WIN [47], W2CB [49], and NOTWB [45]. As a real-world applications experiment, we compare the algorithms in the subset posterior aggregation setting, a standard experiment in the barycentre literature [85, 54, 28, 49]. The barycentre of subset posteriors is known to be close to the true posterior [84], and as such it can be efficient to compute posteriors on only subsets of datasets before aggregating them using a barycentre algorithm. We consider the experimental setup and dataset used in [49] (the same dataset was also used previously in [54, 28]), which uses Poisson and negative-binomial regressions on a bike-rental dataset [29], and report results for the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric. We ran TreeDSBM for 4 IMF iterations with $\varepsilon=0.001$; for full experimental details, see Appendix D.3.
From Table 3, we see that all methods perform strongly. Given that we do not have perfect access to the ground truth barycentre, it is difficult to conclusively say which performs best, but we see that TreeDSBM certainly performs competitively with these state-of-the-art approaches. Moreover, we found TreeDSBM to display fast training—both TreeDSBM and NOTWB obtained good results after only around 3 minutes of training, while W2CB took approximately 10 minutes and WIN took around 45 minutes. For more details regarding runtimes, see Appendix D.3.
Higher-dimensional Gaussian experiments
We also report results for computing the barycentre of Gaussian distributions, in increasingly high dimensions. This is a standard experiment in the literature, because in this setting the ground-truth barycentre is also Gaussian and the parameters can be calculated accurately using a fixed point method [4].
We follow the experimental setup previously used in [49, 47, 45], in which 3 Gaussian distributions and its ground-truth ($\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3}$)-barycentre are randomly generated, for each dimension in $\{64,96,128\}$. We report results for the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ and $\mathrm{L}^{2}\mathrm{-UVP}$ metrics. We see that all the methods again perform well in this setting. For the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric, W2CB and NOTWB appear to have a slight edge, but TreeDSBM is only slightly higher and is comparable with results for WIN. For the $L^{2}\mathrm{-UVP}$ metric, results for TreeDSBM are higher than for W2CB and NOTWB, though the results are still low and are again comparable with WIN.
Table 3: $\mathrm{BW}_{2}^{2}\mathrm{-UVP},\%$ for different algorithms on the subset posterior aggregation experiment in [49], evaluated using 100,000 samples.
@@
$\downarrow$ Negative Binomial 0.009 0.024 0.018 0.012
@@
$d=64$ 0.20 0.04 0.08 0.14
$\downarrow\mathrm{BW}_{2}^{2}\mathrm{-UVP},\%$ $d=96$ 0.30 0.07 0.10 0.15
$d=128$ 0.38 0.12 0.14 0.27
$d=64$ 0.96 0.17 0.10 1.18
$\downarrow L^{2}\mathrm{-UVP},\%$ $d=96$ 1.20 0.20 0.10 1.13